OpenStudy (anonymous):

Tell whether the series converges or diverges. If it converges, give its sum

OpenStudy (tkhunny):

Diverges. Oh, wait...

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}\sin ^{n}(\frac{ \pi }{ 4 }+npi)\]

OpenStudy (anonymous):

LOL

OpenStudy (tkhunny):

What do you glean from the first few terms? Might it be helpful to expand the sine of a sum?

OpenStudy (anonymous):

Ummm

OpenStudy (anonymous):

Well the first term is 1

OpenStudy (anonymous):

The second is root 2/2

OpenStudy (anonymous):

negative root2/2

OpenStudy (tkhunny):

Okay, that's enough of that. Remember your trig formula for the sine for a sum? Time to break it out. \(\sin(\pi/4 + n\pi) = \)

OpenStudy (anonymous):

Do you mean the formula for a sine series?

OpenStudy (tkhunny):

No. Just trigonometry. \(\sin(\alpha+\beta) = \sin(\alpha)\cos(\beta)+\cos(\alpha)\sin(\beta)\) Ringing bells? You're supposed to remember this stuff. :-)

OpenStudy (anonymous):

Nope. I do remember the book giving us that formula for a later problem though

OpenStudy (anonymous):

Actually I do vaguely recall using that formula

OpenStudy (tkhunny):

Well, there you go. What do you get with that expansion? Quiet Voice in Your Head: "Remember your trig..."

OpenStudy (anonymous):

sin(pi/4)cos(npi)+sin(npi)cos(pi/4) So (root2/2)cos(npi)+root2/2(sin(npi))

OpenStudy (anonymous):

That's probably not what you're looking for though

OpenStudy (tkhunny):

No, that's perfect. Now, let's think about it. \(\sin(n\pi) = 0\) for all integers, n, right? We can just throw out that term. What say you of the other term.

OpenStudy (anonymous):

Yes. And cos(npi) oscillates between -1 and 1 for integers

OpenStudy (tkhunny):

Super. So the actual nth term is \(\left(\dfrac{(-1)^{n-1}}{\sqrt{2}}\right)^{n}\). Can we conclude anything about that?

OpenStudy (anonymous):

How did you get that so quickly?

OpenStudy (anonymous):

Oh wait

OpenStudy (anonymous):

You just did the a(r)^(n-1) thing and raised it to the nth power because it is the sine of the nth power

OpenStudy (anonymous):

I see

OpenStudy (tkhunny):

So, now what? Is it an alternating series, or not. \(\left((-1)^{n-1}\right)^{n}\) What can you say about the resolved exponent on -1?

OpenStudy (anonymous):

Wait but it is root 2 over 2

OpenStudy (anonymous):

Not just root 2

OpenStudy (tkhunny):

Right, but let's just worry about one piece at a time. This part is easy: \(\left(\dfrac{1}{\sqrt{2}}\right)^{n}\)

OpenStudy (anonymous):

The series alternates between -root2/2 and root2/2

OpenStudy (tkhunny):

\(\dfrac{\sqrt{2}}{2} = \dfrac{1}{\sqrt{2}}\)!

OpenStudy (tkhunny):

Are you SURE it alternates? This is my question.

OpenStudy (anonymous):

I'm pretty sure

OpenStudy (anonymous):

(root2/2 cos(npi))^n

OpenStudy (anonymous):

and cos(npi) alternates

OpenStudy (anonymous):

Wat

OpenStudy (tkhunny):

Sorry about that... n * (n-1) is always odd. Can you prove it? If n * (n-1) is always odd, then it is NOT alternating.

OpenStudy (anonymous):

I still dont understand. Where did you get n(n-1)

OpenStudy (tkhunny):

It's the exponent on -1.

OpenStudy (anonymous):

There's no -1 in (root2/2 cos(npi))^n

OpenStudy (tkhunny):

There was a minute ago. You lost it along the way. If n is even, then \(\cos(n\pi) = 1\). If n is odd, then \(\cos(n\pi) = -1\). This is exactly equivalent to \((-1)^{n-1}\)

OpenStudy (anonymous):

There's no way my teacher would expect us to recognize that

OpenStudy (tkhunny):

That's why you were given the trig expansion. You already noticed it on your own when you said, "The series alternates between -root2/2 and root2/2".

OpenStudy (anonymous):

Yes. I amend that statement to it alternates between (-root2/2)^n and (root2/2)^n

OpenStudy (tkhunny):

The fact of the matter is, \(\sin(\pi/4 + n\pi) = \left(\dfrac{(-1)^{n-1}}{\sqrt{2}}\right)^{n}\). This is what you just said, agreed?

OpenStudy (tkhunny):

... Sorry, with the n-exponent on the sine. Typo.

OpenStudy (anonymous):

I guess. I dont know where the 2 went

OpenStudy (anonymous):

And I dont know where the (-1)^n-1 came from

OpenStudy (tkhunny):

What 2? It's arithmetic. \(\dfrac{\sqrt{2}}{2} = \dfrac{1}{\sqrt{2}}\)

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Sorry my brain is fried

OpenStudy (tkhunny):

How did you determine that it was alternating?

OpenStudy (anonymous):

Well cosine alternaes

OpenStudy (tkhunny):

Amend: How did you CORRECTLY determine that it was alternating?

OpenStudy (tkhunny):

...and that is all I have written. If n is an integer, then \(\cos(n\pi) = (-1)^{n-1}\). Agreed?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I get that now

OpenStudy (tkhunny):

So, now we have to buy the whole structure. \(\sin(\pi/4 + n\pi) = \dfrac{(-1)^{n-1}}{\sqrt{2}}\) \(\huge{IF}\) n is an integer, agreed? I have deliberately ignored the big exponent at this point. We'll put it back in a minute.

OpenStudy (anonymous):

Ok. Is this if n=0 or if n=1? n=1 right?

OpenStudy (tkhunny):

Certainly for 0 and 1, but also ANY integer. Make sure you believe it.

OpenStudy (anonymous):

I meant for sigma notation

OpenStudy (anonymous):

Its from n=1 to infinity right?

OpenStudy (tkhunny):

That's a fair question, but actually no. Let's fix your fears about that. Is this \((-1)^{n-1}\) ANY differetn from \((-1)^{n+1}\) ?

OpenStudy (tkhunny):

* different

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

So, if we use +1 instead of -1, do you feel better about leaving it at n = 0 for starters?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Because its weird to see the n-1 when you are plugging in 0

OpenStudy (tkhunny):

Fair enough. Now, the ENTIRE argument. \(\sin^{n}(\pi/4 + n\pi) = \left(\dfrac{(-1)^{n+1}}{\sqrt{2}}\right)^{n}\). Do you believe?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

How about this? \(\sin^{n}(\pi/4 + n\pi) = \dfrac{\left((-1)^{n+1}\right)^{n}}{\sqrt{2}^{n}}\).

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

Okay, so the total exponent in the numerator is n*(n+1). Is n*(n+1) Odd or Even - Always, Sometimes, or Never? Think about the parity of each factor.

OpenStudy (anonymous):

OOOH

OpenStudy (tkhunny):

If n is even, then n+1 is what? If n is odd, then n+1 is what?

OpenStudy (anonymous):

If n is even, n+1 is odd If n is odd, n+1 is even

OpenStudy (tkhunny):

So, they NEVER match up. There is always one odd and one even. Agreed?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

What does that mean for the product?

OpenStudy (anonymous):

If n is even, the product will be even If n is odd, the product will be odd

OpenStudy (anonymous):

Wait no

OpenStudy (anonymous):

If n is odd, the product will be even

OpenStudy (anonymous):

Wait I'm wrong again. I go with my original statement

OpenStudy (tkhunny):

Carefully... If n is even, n+1 is odd and n*(n+1) is even*odd and that is always odd. If n is odd, n+1 is even and n*(n+1) is odd*even and that is always odd. That product is ALWAYS odd. You need to believe this.

OpenStudy (anonymous):

The second part is incorrect

OpenStudy (anonymous):

3x2=6

OpenStudy (anonymous):

Thats even

OpenStudy (tkhunny):

So, is it ALWAYS even?

OpenStudy (anonymous):

I'm talking about the second part of your statement where you said an odd times an even is odd

OpenStudy (tkhunny):

How about the first part where I said even * odd is odd. Is that also incorrect?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

Wait no

OpenStudy (anonymous):

Thats correct

OpenStudy (anonymous):

Wait no its not

OpenStudy (anonymous):

2x3=6

OpenStudy (anonymous):

WHAT AM I THINKING

OpenStudy (tkhunny):

You have it! Don't give up! Correct, now! If n is even, n+1 is odd and n*(n+1) is even*odd and that is always even. If n is odd, n+1 is even and n*(n+1) is odd*even and that is always even. That product is ALWAYS even You need to believe this.

OpenStudy (anonymous):

Ok

OpenStudy (tkhunny):

One last thing on this. \(\sin^{n}(\pi/4 + n\pi) = \dfrac{1}{\sqrt{2}^{n}}\) Do you believe?

OpenStudy (anonymous):

I think so?